OpenStudy (anonymous):

in integration, if the derivative of a trigonometric function is given, what to do then ? add 1 in the power of that function and divide ? for example sinx.cosx dx ....

OpenStudy (anonymous):

A contractor employed 210 men to build a house in 60 days. After 12 days , he was joined by 70 more men. In how many days will the remaining work be finished.

OpenStudy (anonymous):

A contractor employed 210 men to build a house in 60 days. After 12 days , he was joined by 70 more men. In how many days will the remaining work be finished.

OpenStudy (unklerhaukus):

use \(\cos(\theta)\sin(\phi)=\frac12\left[\sin(\theta+\phi)-\sin(\theta-\phi)\right]\)